hapter 9 Suppose youll have an annual nominal income of S40,000 for inflation rate is 5 percent per year a. Find the real value of your $40,000 salary for each of the next 3 years. each of the next 3 years, and the Suppose you have a COLA (Cost of Living Adjustment) of 5 percent per year in your contract, which raises your $40,000 salary by 5 percent for each of the next 3 years. Given the 5 percent inflation rate for each of those 3 years, what is the real value of your salary for each year? b.
Answer #1

A)

Real value in year 1 = 40,000(1-5%) = $38,000

Real value in year 2 = 40,000(1-5%)2 = $36,100

Real value in year 3 = 40,000(1-5%)3 = $34,295

[Formula: Real value = Nominal value(1-inflation)]

B)

Cost of living adjustment = 5% = inflation rate

Thus, real value post inflation and COLA will remain the same each year

Thus,

Income in year 1 = Income in year 2 = Income in year 3 = $40,000

[Formula: Real value = Nominal value(1-inflation+COLA)]